# TUTORIAL ARTICLE Contraction Theory for Nonlinear Stability Analysis and Learning-based Control: A Tutorial Overview

### v7

@@ 6.2 NCMs as Lyapunov Functions
where $A$ is the SDC matrix of the system (6.1) given by Lemma 3.1, i.e., $A(\varrho,x,x_{d},u_{d},t)(x-x_{d})=f(x,t)+B(x,t)u_{d}-f(x_{d},t)-B(x_{d},t)u_{d}$. Using the result of Theorem 2.3, one of the Lyapunov functions for the virtual system (6.28) with unknown $K$ can be given as the following transformed squared length integrated over $x$ of (6.1)/(6.2) and $x_{d}$ of (6.3):
@@
where $\mathcal{M}$ defines the NCM of Definition 6.2 modeling $M$ of the CV-STEM contraction metric in Theorem 4.2.
@@
Suppose (6.1) and (6.2) are controlled by $u=u_{d}-K^{*}(x,x_{d},u_{d},t)\mathtt{e}$, where $\mathtt{e}=x-x_{d}$, and that $K^{*}$ is designed by the following convex optimization for given $(x,x_{d},u_{d},t)$:
@@
where $\mathcal{M}$ defines the NCM of Definition 6.2 that models $M$ of the CV-STEM contraction metric in Theorem 4.2 as in (6.22), $\mathcal{C}_{K}$ is a convex set containing admissible $K$, $\alpha$ is the contraction rate, and $\beta$ is as defined in theorem 3.1 (i.e., $\beta=0$ for deterministic systems and $\beta=\alpha_{s}$ for stochastic systems). Then Theorems 2.4 and 2.5 still hold for the Lyapunov function defined in (6.29), yielding the following bounds:
@@
where $\bar{p}=\sup_{x,x_{d},u_{d},t}p$, $\underline{m}\mathrm{I}\preceq\mathcal{M}\preceq\overline{m}\mathrm{I}$ as in (2.32), $\alpha_{d}\in\mathbb{R}_{>0}$ is an arbitrary positive constant (see (6.35)) selected to satisfy $\alpha_{p}=\alpha-\alpha_{d}/2>0$, $C_{C}=\bar{g}_{c}^{2}({2}{\alpha_{G}}^{-1}+1)$, $\alpha_{G}\in\mathbb{R}_{>0}$ is an arbitrary constant as in Theorem 2.5, $\varepsilon\in\mathbb{R}_{\geq 0}$, and the disturbance terms $\bar{d}_{c}$ and $\bar{g}_{c}$ are given in (6.1) and (6.2), respectively. Note that the problem (6.30) is feasible due to the constraint relaxation variable $p$.
@@
which gives (6.33) and (6.34) due to Lemma 2.2. If $\epsilon_{\ell}=0$ in (6.22), $K=R^{-1}B^{\top}M$ satisfies (6.31) with $p=0$ as we have $\mathcal{M}=M$ and $M$ defines the contraction metric of Theorem 4.2, which implies that (6.30) with $p=0$ is always feasible. $\qed$
We could also use the CCM-based differential feedback formulation of Theorem 4.6 in Theorem 6.4 (see Manchester and Slotine [2017], Singh et al. [2017]). To this end, define $E_{\mathrm{NCM}}$ (Riemannian energy) as follows:
@@
where $\mathcal{M}$ defines the NCM of Definition 6.2 that models $M$ of the CCM in Theorem 4.6, $\gamma$ is the minimizing geodesic connecting $x(t)=\gamma(1,t)$ of (6.1) and $x_{d}(t)=\gamma(0,t)$ of (6.3) (see Theorem 2.3), and $\gamma_{\mu}={\partial\gamma}/{\partial\mu}$.
@@
Suppose $u=u^{*}$ of (6.1) is designed by the following convex optimization for given $(x,x_{d},u_{d},t)$:
@@
s.t. $\begin{aligned} &\text{{\rm s.t.}{} }\frac{\partial E}{\partial t}+2\gamma_{\mu}(1,t)^{\top}\mathcal{M}(x,t)(f(x,t)+B(x,t)u)\\ &\text{{\color[rgb]{1,1,1}{\rm s.t.}{}} }-2\gamma_{\mu}(0,t)^{\top}\mathcal{M}(x_{d},t)(f(x_{d},t)+B(x_{d},t)u_{d})\leq-2\alpha E+p\end{aligned}$ (6.39)
where $\mathcal{M}$ defines the NCM of Definition 6.2 that models the CCM of Theorem 4.6, and $\mathcal{C}_{u}$ is a convex set containing admissible $u$. Then the bound (6.32) of Theorem 6.4 holds.
@@
Since the right-hand side of (6.39) represents $\dot{E}$ of (6.37) if $d_{c}=0$ in (6.1), the comparison lemma of Lemma 2.1 gives (6.32) as in Theorem 6.4. The rest follows from the fact that $u$ of Theorem 4.6 is a feasible solution of (6.38) if $p=0$ Manchester and Slotine [2017], Singh et al. [2017].
@@
As discussed in Remark 4.2, Theorem 6.5 can also be formulated in a stochastic setting as in Theorem 6.4 using the technique discussed in Tsukamoto et al. [2021].
Theorems 6.4 and 6.5 provide a perspective on the NCM stability different from Theorems 6.1 and 6.3 written in terms of the constraint relaxation variable $p$, by directly using the NCM as a contraction metric instead of the CV-STEM contraction metric or the CCM. Since the CLF problems are feasible with $p=0$ when we have zero NCM modeling error, it is implied that the upper bound of $p$ (i.e., $\bar{p}$ in (6.32) and (6.33)) decreases as we achieve a smaller learning error $\epsilon_{\ell}$ using more training data for verifying (6.22) (see Remark 5.1). This intuition will be formalized later in Theorem 9.3 of Sec. 9 Boffi et al. [2020]. See Manchester and Slotine [2017], Singh et al. [2017] for how we compute the minimizing geodesic $\gamma$ and its associated Riemannian energy $E$ in Theorem 6.5.
@@
If $\mathcal{C}_{K}=\mathbb{R}^{m\times n}$ and $\mathcal{C}_{u}=\mathbb{R}^{m}$ in (6.30) and (6.38), respectively, they can both be expressed as the following quadratic optimization problem:
@@
by defining ${c}\in\mathbb{R}^{\mathtt{v}}$, $\varphi_{0}\in\mathbb{R}$, $\varphi_{1}\in\mathbb{R}^{\mathtt{v}}$, and $v\in\mathbb{R}^{\mathtt{v}}$ appropriately. Applying the KKT condition [Boyd and Vandenberghe, 2004, pp. 243-244] to this problem, we can show that
@@
When $q$ is parameterized linearly by $\mu$ in Theorem 6.4, i.e., $q=\mu x+(1-\mu)x_{d}$, we have $V_{\mathrm{NCM}}=(x-x_{d})^{\top}\mathcal{M}(x-x_{d})$ as in the standard Lyapunov function formulation Sontag [1983, 1989], Khalil [2002]. Also, a linear input constraint, $u_{\min}\leq u\leq u_{\max}$ can be implemented by using
@@
For the CV-STEM and NCM state estimation in Theorems 4.5, 6.2, and 6.3, we could design a similar Lyapunov function-based estimation scheme if its contraction constraints of Theorem 4.5 do not explicitly depend on the actual state $x$. This can be achieved by using $A(\varrho_{A},\hat{x},t)$ and $C(\varrho_{C},\hat{x},t)$ instead of $A(\varrho_{A},x,\hat{x},t)$ and $C(\varrho_{C},x,\hat{x},t)$ in Theorem 4.5 as in Theorem 3.2 Dani et al. [2015], leading to exponential boundedness of system trajectories as derived in Tsukamoto and Chung [2021], Dani et al. [2015], Tsukamoto and Chung [2019].
6.3 Remarks in NCM Implementation
This section briefly summarizes several practical techniques in constructing NCMs using the CV-STEM of Theorems 4.24.5, and 4.6.
6.3.1 Neural Network Training
We model $M$ of the CV-STEM contraction metric sampled offline by the DNN of Definition 6.1, optimizing its hyperparameters by stochastic gradient descent Robbins and Monro [1951], Wensing and Slotine [2020] to satisfy the learning error condition (5.3) (see Remark 5.1). The positive definiteness of $M$ could be exploited to reduce the dimension of the DNN target output due to the following lemma Tsukamoto and Chung [2021].
@@ 6.3.1 Neural Network Training
See [Horn and Johnson, 2012, pp. 441].
Selecting $\Theta$ of $M=\Theta^{\top}\Theta$ as the unique Cholesky decomposition of $M$ and training the DNN using only the non-zero entries of $\Theta$, we can reduce the dimension of the target output by n(n1)/2n(n-1)/2 without losing any information on $M$. The pseudocode to obtain NCMs, using this approach depicted in Fig. 3, can be found in Tsukamoto and Chung [2021].
Instead of solving the convex optimization in Theorems 4.24.5, and 4.6 to sample training data, we could also use them directly for training and optimizing the DNN hyperparameters, treating the constraints and the objective functions as the DNN loss functions as demonstrated in Sun et al. [2020]. Although this approach no longer preserves the convexity and can lead only to a sub-optimal solution, this still gives the exponential tracking error bounds as long as we can get sufficiently small $\epsilon_{\ell 0}$ and $\epsilon_{\ell 1}$ in the learning error assumption (5.3), as discussed in Theorems 5.25.3, and 6.16.3. See Tsukamoto et al. [2021] for more details on the robustness and stability properties of this approach.
6.3.2 Lipschitz Condition and Spectral Normalization
Let us first define a spectrally-normalized DNN, a useful mathematical tool designed to overcome the instability of network training by constraining (6.6) of Definition 6.1 to be Lipschitz, i.e., $\exists\ L_{nn}\in\mathbb{R}_{\geq 0}$ s.t. $\|\varphi(x)-\varphi(x^{\prime})\|\leq L_{nn}\|x-x^{\prime}\|,~\forall x,x^{\prime}$ Miyato et al. [2018], Shi et al. [2019].
@@ 6.3.2 Lipschitz Condition and Spectral Normalization
A spectrally-normalized DNN is a DNN of Definition 6.1 with its weights $W_{\ell}$ normalized as $W_{\ell}=(C_{nn}\Omega_{\ell})/\|\Omega_{\ell}\|$, where $C_{nn}\in\mathbb{R}_{\geq 0}$ is a given constant.
@@
A spectrally-normalized DNN given in Definition 6.3 is Lipschitz continuous with its 2-norm Lipschitz constant $C_{nn}^{L+1}L_{\sigma}^{L}$, where $L_{\sigma}$ is the Lipschitz constant of the activation function $\sigma$ in (6.6). Also, it is robust to perturbation in its input.
@@
Let $\|f\|_{\mathrm{Lip}}$ represent the Lipschitz constant of a function $f$ and let $W_{\ell}=(C_{nn}\Omega_{\ell})/\|\Omega_{\ell}\|$. Using the property $\|f_{1}\circ f_{2}\|_{\mathrm{Lip}}\leq\|f_{1}\|_{\mathrm{Lip}}\|f_{2}\|_{\mathrm{Lip}}$, we have for a spectrally-normalized DNN in Definition 6.3 that
@@
In general, a spectrally-normalized DNN is useful for improving the robustness and generalization properties of DNNs, and for obtaining the learning error bound as in (5.3), as delineated in Remark 5.1 Miyato et al. [2018], Bartlett et al. [2017], Neyshabur et al. [2018]. Using it for system identification yields a DNN-based nonlinear feedback controller with a formal stability guarantee as shall be seen in Sec. 9.1.
For the NCM framework with stochastic perturbation, we can utilize the spectrally-normalized DNN of Definition 6.3 to guarantee the Lipschitz condition on $\partial M/\partial x_{i}$, appeared in the stochastic contraction condition of Theorem 2.5 (see Proposition 1 of Tsukamoto et al. [2021]). We could also utilize the technique proposed in Revay et al. [2020, 2021], Manchester et al. [2021], which designs Lipschitz-bounded equilibrium neural networks using contraction theory, for obtaining a result analogous to Lemma 6.2. The pseudocode for the NCM construction in a stochastic setting can be found in Tsukamoto et al. [2021].
@@
Table 4: Notations in Example 6.5.
@@
Let us demonstrate how having the Lipschitz condition on $\partial M/\partial x_{i}$ in Theorem 2.5 would affect the NCM-based control and estimation performance. We consider a rocket autopilot problem perturbed by stochastic disturbances as in Tsukamoto et al. [2021], the unperturbed dynamics of which is depicted in Fig. 5 and given as follows, assuming that the pitch rate $q$ and specific normal force are available as a measurement via rate gyros and accelerometers Shamma and Cloutier [1993], Bennani et al. [1997]:
@@
where $C(M)=p_{0}M^{2}S$, $\phi_{z}(\varphi,M)=0.7(a_{n}\varphi^{3}+b_{n}\varphi|\varphi|+c_{n}(2+M/3)\varphi)$, $\phi_{m}(\varphi,M)=0.7(a_{m}\varphi^{3}+b_{m}\varphi|\varphi|-c_{m}(7-8M/3)\varphi)$, $\bar{d}_{n}=0.7d_{n}$, $\bar{d}_{m}=0.7d_{m}$, $(a_{n},b_{n},c_{n},d_{n})$ and $(a_{m},b_{m},c_{m},d_{m})$ are given in Bennani et al. [1997], and the notations are defined in Table 4. Since this example explicitly takes into account stochastic perturbation, the spectrally-normalized DNN of Definition 6.3 is used to guarantee the Lipschitz condition of Theorem 2.5 by Lemma 6.2.
Figure 6 shows the state estimation and tracking error performance of each estimator and controller, where the NSCM is the stochastic counterpart of the NCM aforementioned in Remark 6.1. It is demonstrated that the steady-state errors of the NSCM and CV-STEM are indeed smaller than its steady-state upper bounds of Theorems 4.2 and 4.5 (solid black line), while the other controllers violate this condition. In particular, the optimal contraction rate of the deterministic NCM for state estimation is turned out to be much larger than the NSCM as it does not account for stochastic perturbation, which makes the NCM trajectory diverge around $t=5.8$ in Fig. 6. The NSCM circumvents this difficulty by imposing the Lipschitz condition on $\partial M/\partial x_{i}$ as in Theorem 2.5 using spectral normalization of Lemma 6.2. See Tsukamoto et al. [2021] for additional details.
Figure 6: Rocket state estimation and tracking errors in Example 6.5 ($x=[\varphi,q]^{T}$).
7 Learning-based Robust Motion Planning

### v8

@@ 6.2. NCMs as Lyapunov Functions
where $A$ is the SDC matrix of the system (6.1) given by Lemma 3.1, i.e., $A(\varrho,x,x_{d},u_{d},t)(x-x_{d})=f(x,t)+B(x,t)u_{d}-f(x_{d},t)-B(x_{d},t)u_{d}$. Using the result of Theorem LABEL:THM:Thm:path_integral, one of the Lyapunov functions for the virtual system (6.28) with unknown $K$ can be given as the following transformed squared length integrated over $x$ of (6.1)/(6.2) and $x_{d}$ of (6.3):
@@
where $\mathcal{M}$ defines the NCM of Definition LABEL:DEF:Def:NCM modeling $M$ of the CV-STEM contraction metric in Theorem LABEL:THM:Thm:CV-STEM.
@@
Control Lyapunov Functions with Neural Contraction MetricsThm:ncm_clf Suppose (6.1) and (6.2) are controlled by $u=u_{d}-K^{*}(x,x_{d},u_{d},t)\mathtt{e}$, where $\mathtt{e}=x-x_{d}$, and that $K^{*}$ is designed by the following convex optimization for given $(x,x_{d},u_{d},t)$:
@@
where $\mathcal{M}$ defines the NCM of Definition LABEL:DEF:Def:NCM that models $M$ of the CV-STEM contraction metric in Theorem LABEL:THM:Thm:CV-STEM as in (6.22), $\mathcal{C}_{K}$ is a convex set containing admissible $K$, $\alpha$ is the contraction rate, and $\beta$ is as defined in theorem LABEL:THM:Thm:CVSTEM:LMI (i.e., $\beta=0$ for deterministic systems and $\beta=\alpha_{s}$ for stochastic systems). Then Theorems LABEL:THM:Thm:Robust_contraction_original and LABEL:THM:Thm:robuststochastic still hold for the Lyapunov function defined in (6.29), yielding the following bounds:
@@
where $\bar{p}=\sup_{x,x_{d},u_{d},t}p$, $\underline{m}\mathrm{I}\preceq\mathcal{M}\preceq\overline{m}\mathrm{I}$ as in (2.32), $\alpha_{d}\in\mathbb{R}_{>0}$ is an arbitrary positive constant (see (6.35)) selected to satisfy $\alpha_{p}=\alpha-\alpha_{d}/2>0$, $C_{C}=\bar{g}_{c}^{2}({2}{\alpha_{G}}^{-1}+1)$, $\alpha_{G}\in\mathbb{R}_{>0}$ is an arbitrary constant as in Theorem LABEL:THM:Thm:robuststochastic, $\varepsilon\in\mathbb{R}_{\geq 0}$, and the disturbance terms $\bar{d}_{c}$ and $\bar{g}_{c}$ are given in (6.1) and (6.2), respectively. Note that the problem (6.30) is feasible due to the constraint relaxation variable $p$.
@@
which gives (6.33) and (6.34) due to Lemma 2.2. If $\epsilon_{\ell}=0$ in (6.22), $K=R^{-1}B^{\top}M$ satisfies (6.31) with $p=0$ as we have $\mathcal{M}=M$ and $M$ defines the contraction metric of Theorem LABEL:THM:Thm:CV-STEM, which implies that (6.30) with $p=0$ is always feasible. $\qed$
We could also use the CCM-based differential feedback formulation of Theorem LABEL:THM:Thm:ccm_cvstem in Theorem LABEL:THM:Thm:ncm_clf (see ccm, 7989693). To this end, define $E_{\mathrm{NCM}}$ (Riemannian energy) as follows:
@@
where $\mathcal{M}$ defines the NCM of Definition LABEL:DEF:Def:NCM that models $M$ of the CCM in Theorem LABEL:THM:Thm:ccm_cvstem, $\gamma$ is the minimizing geodesic connecting $x(t)=\gamma(1,t)$ of (6.1) and $x_{d}(t)=\gamma(0,t)$ of (6.3) (see Theorem LABEL:THM:Thm:path_integral), and $\gamma_{\mu}={\partial\gamma}/{\partial\mu}$.
@@
Control Lyapunov Functions with Control Contraction MetricsThm:ncm_ccm_clf Suppose $u=u^{*}$ of (6.1) is designed by the following convex optimization for given $(x,x_{d},u_{d},t)$:
@@
s.t. $\begin{aligned} &\text{{\rm s.t.}{} }\frac{\partial E}{\partial t}+2\gamma_{\mu}(1,t)^{\top}\mathcal{M}(x,t)(f(x,t)+B(x,t)u)\\ &\text{{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}{\rm s.t.}{}} }-2\gamma_{\mu}(0,t)^{\top}\mathcal{M}(x_{d},t)(f(x_{d},t)+B(x_{d},t)u_{d})\leq-2\alpha E+p\end{aligned}$ (6.39)
where $\mathcal{M}$ defines the NCM of Definition LABEL:DEF:Def:NCM that models the CCM of Theorem LABEL:THM:Thm:ccm_cvstem, and $\mathcal{C}_{u}$ is a convex set containing admissible $u$. Then the bound (6.32) of Theorem LABEL:THM:Thm:ncm_clf holds.
@@
Since the right-hand side of (6.39) represents $\dot{E}$ of (6.37) if $d_{c}=0$ in (6.1), the comparison lemma of Lemma 2.1 gives (6.32) as in Theorem LABEL:THM:Thm:ncm_clf. The rest follows from the fact that $u$ of Theorem LABEL:THM:Thm:ccm_cvstem is a feasible solution of (6.38) if $p=0$ ccm, 7989693.
@@
As discussed in Remark 4.2, Theorem LABEL:THM:Thm:ncm_ccm_clf can also be formulated in a stochastic setting as in Theorem LABEL:THM:Thm:ncm_clf using the technique discussed in cdc_ncm.
Theorems LABEL:THM:Thm:ncm_clf and LABEL:THM:Thm:ncm_ccm_clf provide a perspective on the NCM stability different from Theorems LABEL:THM:Thm:NCMstability1 and LABEL:THM:Thm:NCMstability2 written in terms of the constraint relaxation variable $p$, by directly using the NCM as a contraction metric instead of the CV-STEM contraction metric or the CCM. Since the CLF problems are feasible with $p=0$ when we have zero NCM modeling error, it is implied that the upper bound of $p$ (i.e., $\bar{p}$ in (6.32) and (6.33)) decreases as we achieve a smaller learning error $\epsilon_{\ell}$ using more training data for verifying (6.22) (see Remark 5.1). This intuition will be formalized later in Theorem LABEL:THM:Thm:data_driven_cont_thm of Sec. 9 boffi2020learning. See ccm, 7989693 for how we compute the minimizing geodesic $\gamma$ and its associated Riemannian energy $E$ in Theorem LABEL:THM:Thm:ncm_ccm_clf.
@@
Analytical Solutions of CLF Convex Optimization If $\mathcal{C}_{K}=\mathbb{R}^{m\times n}$ and $\mathcal{C}_{u}=\mathbb{R}^{m}$ in (6.30) and (6.38), respectively, they can both be expressed as the following quadratic optimization problem:
@@
by defining ${c}\in\mathbb{R}^{\mathtt{v}}$, $\varphi_{0}\in\mathbb{R}$, $\varphi_{1}\in\mathbb{R}^{\mathtt{v}}$, and $v\in\mathbb{R}^{\mathtt{v}}$ appropriately. Applying the KKT condition [citeulike:163662, pp. 243-244] to this problem, we can show that
@@
Control Input Constraints When $q$ is parameterized linearly by $\mu$ in Theorem LABEL:THM:Thm:ncm_clf, i.e., $q=\mu x+(1-\mu)x_{d}$, we have $V_{\mathrm{NCM}}=(x-x_{d})^{\top}\mathcal{M}(x-x_{d})$ as in the standard Lyapunov function formulation doi:10.1137/0321028, SONTAG1989117, Khalil:1173048. Also, a linear input constraint, $u_{\min}\leq u\leq u_{\max}$ can be implemented by using
@@
For the CV-STEM and NCM state estimation in Theorems LABEL:THM:Thm:CV-STEM-estimation, LABEL:THM:Thm:NCMstability1_est, and LABEL:THM:Thm:NCMstability2, we could design a similar Lyapunov function-based estimation scheme if its contraction constraints of Theorem LABEL:THM:Thm:CV-STEM-estimation do not explicitly depend on the actual state $x$. This can be achieved by using $A(\varrho_{A},\hat{x},t)$ and $C(\varrho_{C},\hat{x},t)$ instead of $A(\varrho_{A},x,\hat{x},t)$ and $C(\varrho_{C},x,\hat{x},t)$ in Theorem LABEL:THM:Thm:CV-STEM-estimation as in Theorem LABEL:THM:Thm:fixed_sdc Ref:Stochastic, leading to exponential boundedness of system trajectories as derived in mypaperTAC, Ref:Stochastic, mypaper.
6.3. Remarks in NCM Implementation
This section briefly summarizes several practical techniques in constructing NCMs using the CV-STEM of Theorems $LABEL:THM:Thm:CV-STEM$-estimation, and LABEL:THM:Thm:ccm_cvstem.
6.3.1. Neural Network Training
We model $M$ of the CV-STEM contraction metric sampled offline by the DNN of Definition LABEL:DEF:Def:NN_definition, optimizing its hyperparameters by stochastic gradient descent sgd, beyondconvexity to satisfy the learning error condition (5.3) (see Remark 5.1). The positive definiteness of $M$ could be exploited to reduce the dimension of the DNN target output due to the following lemma ncm.
@@ 6.3.1. Neural Network Training
See [10.5555/2422911, pp. 441].
Selecting $\Theta$ of $M=\Theta^{\top}\Theta$ as the unique Cholesky decomposition of $M$ and training the DNN using only the non-zero entries of $\Theta$, we can reduce the dimension of the target output by n(n1)/2n(n-1)/2 without losing any information on $M$. The pseudocode to obtain NCMs, using this approach depicted in Fig. 3, can be found in ncm.
Instead of solving the convex optimization in Theorems $LABEL:THM:Thm:CV-STEM$-estimation, and LABEL:THM:Thm:ccm_cvstem to sample training data, we could also use them directly for training and optimizing the DNN hyperparameters, treating the constraints and the objective functions as the DNN loss functions as demonstrated in chuchu. Although this approach no longer preserves the convexity and can lead only to a sub-optimal solution, this still gives the exponential tracking error bounds as long as we can get sufficiently small $\epsilon_{\ell 0}$ and $\epsilon_{\ell 1}$ in the learning error assumption (5.3), as discussed in Theorems LABEL:THM:Thm:contraction_learningLABEL:THM:Thm:contraction_learning_sto, and LABEL:THM:Thm:NCMstability1LABEL:THM:Thm:NCMstability2. See cdc_ncm for more details on the robustness and stability properties of this approach.
6.3.2. Lipschitz Condition and Spectral Normalization
Let us first define a spectrally-normalized DNN, a useful mathematical tool designed to overcome the instability of network training by constraining (6.6) of Definition LABEL:DEF:Def:NN_definition to be Lipschitz, i.e., $\exists\ L_{nn}\in\mathbb{R}_{\geq 0}$ s.t. $\|\varphi(x)-\varphi(x^{\prime})\|\leq L_{nn}\|x-x^{\prime}\|,~\forall x,x^{\prime}$ miyato2018spectral, neurallander.
@@ 6.3.2. Lipschitz Condition and Spectral Normalization
Spectrally-Normalized DNNsDef:SN A spectrally-normalized DNN is a DNN of Definition LABEL:DEF:Def:NN_definition with its weights $W_{\ell}$ normalized as $W_{\ell}=(C_{nn}\Omega_{\ell})/\|\Omega_{\ell}\|$, where $C_{nn}\in\mathbb{R}_{\geq 0}$ is a given constant.
@@
A spectrally-normalized DNN given in Definition LABEL:DEF:Def:SN is Lipschitz continuous with its 2-norm Lipschitz constant $C_{nn}^{L+1}L_{\sigma}^{L}$, where $L_{\sigma}$ is the Lipschitz constant of the activation function $\sigma$ in (6.6). Also, it is robust to perturbation in its input.
@@
Let $\|f\|_{\mathrm{Lip}}$ represent the Lipschitz constant of a function $f$ and let $W_{\ell}=(C_{nn}\Omega_{\ell})/\|\Omega_{\ell}\|$. Using the property $\|f_{1}\circ f_{2}\|_{\mathrm{Lip}}\leq\|f_{1}\|_{\mathrm{Lip}}\|f_{2}\|_{\mathrm{Lip}}$, we have for a spectrally-normalized DNN in Definition LABEL:DEF:Def:SN that
@@
In general, a spectrally-normalized DNN is useful for improving the robustness and generalization properties of DNNs, and for obtaining the learning error bound as in (5.3), as delineated in Remark 5.1 miyato2018spectral, marginbounds1, marginbounds2. Using it for system identification yields a DNN-based nonlinear feedback controller with a formal stability guarantee as shall be seen in Sec. 9.1.
For the NCM framework with stochastic perturbation, we can utilize the spectrally-normalized DNN of Definition LABEL:DEF:Def:SN to guarantee the Lipschitz condition on $\partial M/\partial x_{i}$, appeared in the stochastic contraction condition of Theorem LABEL:THM:Thm:robuststochastic (see Proposition 1 of nscm). We could also utilize the technique proposed in revay2020lipschitz, revay2021recurrent, cdc_systemid, which designs Lipschitz-bounded equilibrium neural networks using contraction theory, for obtaining a result analogous to Lemma 6.2. The pseudocode for the NCM construction in a stochastic setting can be found in nscm.
@@
Table 4: Notations in Example LABEL:EX:ex:rocket.
@@
Rocket Autopilotex:rocket Let us demonstrate how having the Lipschitz condition on $\partial M/\partial x_{i}$ in Theorem LABEL:THM:Thm:robuststochastic would affect the NCM-based control and estimation performance. We consider a rocket autopilot problem perturbed by stochastic disturbances as in nscm, the unperturbed dynamics of which is depicted in Fig. 5 and given as follows, assuming that the pitch rate $q$ and specific normal force are available as a measurement via rate gyros and accelerometers doi:10.2514/3.20997, doi:10.2514/6.1997-3641:
@@
where $C(M)=p_{0}M^{2}S$, $\phi_{z}(\varphi,M)=0.7(a_{n}\varphi^{3}+b_{n}\varphi|\varphi|+c_{n}(2+M/3)\varphi)$, $\phi_{m}(\varphi,M)=0.7(a_{m}\varphi^{3}+b_{m}\varphi|\varphi|-c_{m}(7-8M/3)\varphi)$, $\bar{d}_{n}=0.7d_{n}$, $\bar{d}_{m}=0.7d_{m}$, $(a_{n},b_{n},c_{n},d_{n})$ and $(a_{m},b_{m},c_{m},d_{m})$ are given in doi:10.2514/6.1997-3641, and the notations are defined in Table 4. Since this example explicitly takes into account stochastic perturbation, the spectrally-normalized DNN of Definition LABEL:DEF:Def:SN is used to guarantee the Lipschitz condition of Theorem LABEL:THM:Thm:robuststochastic by Lemma 6.2.
Figure 6 shows the state estimation and tracking error performance of each estimator and controller, where the NSCM is the stochastic counterpart of the NCM aforementioned in Remark 6.1. It is demonstrated that the steady-state errors of the NSCM and CV-STEM are indeed smaller than its steady-state upper bounds of Theorems LABEL:THM:Thm:CV-STEM and LABEL:THM:Thm:CV-STEM-estimation (solid black line), while the other controllers violate this condition. In particular, the optimal contraction rate of the deterministic NCM for state estimation is turned out to be much larger than the NSCM as it does not account for stochastic perturbation, which makes the NCM trajectory diverge around $t=5.8$ in Fig. 6. The NSCM circumvents this difficulty by imposing the Lipschitz condition on $\partial M/\partial x_{i}$ as in Theorem LABEL:THM:Thm:robuststochastic using spectral normalization of Lemma 6.2. See nscm for additional details.
Figure 6: Rocket state estimation and tracking errors in Example LABEL:EX:ex:rocket ($x=[\varphi,q]^{T}$).
7. Learning-based Robust Motion Planning